2-Calculate the payoff at expiration for a put option on the S&P 100 stock index in which the underlying price is 623.22 at expiration, the multiplier is 100, the strike price is a) 475 b) 750
Answer #1

Payoff on put=max(0,Strike price-Stock price)

a)
Payoff=

=max(0,475-623.22)

=0

Hence total payoff shall be zero

b)

=max(0,750-623.22)

=$126.78

Multiplier is 100

Hence total payoff=$126.78*100=$12,678